Find the area of a triangle with base 10 cm and corresponding height 4 cm.
Answer by rajagopalan(174)

base 10 cm
corresponding height 4 cm.
Area = 1/2*base*height
=1/2 * 10 * 4
=5*4
=20 sq cm